OpenStudy (anonymous):

Assume that there are 6 different issues of Time, 7 different issues of Sports Illustrated, and 4 different issues of Popular Science, including the December 1st issue, on the rack. You choose 4 of them at random. What is the probability that you choose the December 1st issue of Popular Science?

OpenStudy (anonymous):

Total number of issues: 18 December 1rst Issue: Only One Popular Science: 4 The chance of picking December 1rst issue is 1/5

OpenStudy (anonymous):

that is wrong 1/5

OpenStudy (anonymous):

We have a total amount of outcomes at: \[ 6+7+4=17 \]Thus, let us compute the chance of NOT picking the Popular Science magazine needed (since each event is independent of one another): \[ P(E_1)P(E_2)P(E_3)=\frac{16}{17}\cdot\frac{15}{16}\cdot\frac{14}{15}\cdot\frac{13}{14}=\\ \frac{_{16}P_4}{_{17}P_4}=\frac{13}{17} \]Therefore the probability of picking the December 1st edition is: \[ 1-P(E_n)=1-\frac{13}{17}=\frac{4}{17} \]And we're done.

Directrix (directrix):

Using combinations, there are C(1,1)]*[C(16,3)] ways to choose the four magazines with the Dec 1 issue being selected [C(1,1)] which leaves C(16,3)] ways to choose the other three magazines. The number of ways to choose 4 magazines from a group of 17 is [C(17,4)]. The probability of selecting 4 magazines with the Dec 1 issue as one of the three is the following: C(1,1)]*[C(16,3)]/[C(17,4)] = 4/17.

OpenStudy (anonymous):

ahh I misread.